OpenStudy (anonymous):

can someone explain to me how to do this. 3x+5<2x-8

OpenStudy (anonymous):

basically solve just like you would an equation 3x<2x-8-5

OpenStudy (anonymous):

3x-2x<-13

OpenStudy (anonymous):

x<-13

OpenStudy (anonymous):

got it? @mimimarie416

OpenStudy (anonymous):

yes thank you so much

OpenStudy (anonymous):

what you want to do is put all the x's on one side. So subtract 2x from both sides. 3x - 2x + 5 = 2x - 2x + 8 On the right side 2x - 2x is 0 so you are left with only 8. On the left side 3x - 2x is only x so now this is your new equation. x + 5 = 8 From here you want to subtract 5 from both sides x + 5 - 5 = 8 - 5 On the left side 5 - 5 is 0 so you are only left with x and on the right side you subtract 8 - 5 which is 3. So the final solution is x = 3. If you want to check to make sure it is right then plug 3 back into the original equation. 3(3) + 5 = 2(3) + 8 9 + 5 = 6 + 8 14 = 14